OpenStudy (ineedhelpnow):

find the quatioiin of the line connecting these two points: (4, 2-) and (-2, 0)

OpenStudy (curry):

is that 4,-2?

OpenStudy (ineedhelpnow):

sry. yes

OpenStudy (curry):

slope = -1/3 y=-1/3x-2/3